# Observations of microlensed images with dual-field interferometry: on-sky demonstration and prospects

### v1

@@ Abstract
Interferometric observations of gravitational microlensing events offer an opportunity for precise, efficient, and direct mass and distance measurements of lensing objects, especially those of isolated neutron stars and black holes. However, such observations were previously possible for only a handful of extremely bright events. The recent development of a dual-field interferometer, GRAVITY Wide, has made it possible to reach out to significantly fainter objects, and increase the pool of microlensing events amenable to interferometric observations by two orders of magnitude. Here, we present the first successful observation of a microlensing event with GRAVITY Wide and the resolution of microlensed images in the event OGLE-2023-BLG-0061/KMT-2023-BLG-0496. We measure the angular Einstein radius of the lens with a sub-percent precision, $\theta_{\rm E}=1.280\pm 0.009$ mas. Combined with the microlensing parallax detected from the event light curve, the mass and distance to the lens are found to be $0.472\pm 0.012\,M_{\odot}$ and $1.81\pm 0.05$ kpc, respectively. We present the procedure for the selection of targets for interferometric observations, and discuss possible systematic effects affecting GRAVITY Wide data. This detection demonstrates the capabilities of the new instrument and it opens up completely new possibilities for the follow-up of microlensing events, and future routine discoveries of isolated neutron stars and black holes.
@@ 1 Introduction
In his seminal paper on gravitational microlensing, Einstein (1936) realized that the angular separation between the two images of the source star created by a Galactic gravitational lens would be well below the resolution limits of contemporary telescopes. That led him to believe that observing the microlensed images would be, in practice, impossible. This notion persisted for many decades until the first optical interferometry facilities were developed.
@@
Although a few tens of thousands of single-lens microlensing events have been discovered so far, only a few had the angular Einstein radius measured. That was possible for only a small number of events exhibiting finite-source effects (Gould, 1994; Nemiroff & Wickramasinghe, 1994; Witt & Mao, 1994), and the probability of such measurements is biased towards low-mass (planetary-mass) lenses. Another route for $\theta_{\rm E}$ measurements involved astrometric observations (Hog et al., 1995; Miyamoto & Yoshii, 1995; Walker, 1995), which led to the first detection of an isolated black hole (Sahu et al., 2022; Lam et al., 2022; Mróz et al., 2022; Lam & Lu, 2023). These observations, however, required long-term astrometric monitoring, which is not feasible for a large sample of events. Conversely, interferometric observations are much more efficient as only one exposure is sufficient to resolve the microlensed images and measure the angular Einstein radius. When combined with the microlensing parallax ($\pi_{\rm E}\equiv\pi_{\rm rel}/\theta_{\rm E}$) measurements from the light curve, they allow us to precisely determine the masses, distances, and transverse velocities of isolated objects, including neutron stars and black holes.
@@
Delplancke et al. (2001) were the first to discuss the prospects of using the European Southern Observatory (ESO) Very Large Telescope Interferometer (VLTI) to study gravitational microlensing events. They estimated that the first-generation VLTI instruments would be able to observe dozens of events every year in the mid-2000s. Yet, the first successful resolution of the microlensed images did not materialize until late 2017, when Dong et al. (2019) observed the microlensing event TCP J05074264+2447555 (a.k.a., Kojima-1) with the second-generation VLTI instrument GRAVITY (GRAVITY Collaboration et al., 2017).
@@
Why did it take almost two decades to achieve this milestone? The early predictions on the VLTI performance underestimated the role and impact of many practical problems encountered during system operations, such as beam stability, vibrations, or air dispersion (e.g., Delplancke, 2008). Moreover, even when these technical problems were solved or mitigated, the performance of a ground-based interferometer is ultimately limited by atmospheric turbulence, which breaks the coherence of wavefronts arriving at each telescope. Therefore, the observations must be carried out with exposure times ($\sim 1$ ms) shorter than the typical wavefront coherence timescale (on the order of 20–30 ms) and shorter than the mechanical vibrations timescale, which dominate in the 10–1000 Hz regime. Even for the largest modern 10-m-class telescopes, this requirement sets a limiting magnitude of $K\approx 10$ for interferometric observations. Gravitational microlensing events brighter than this limit are exceedingly rare (Figure 1). They occur at most a few times a year, and they exceed the GRAVITY limiting magnitude for a short period of time. The three events with the published interferometric observations were unusually bright: TCP J05074264+2447555 (Dong et al., 2019; Zang et al., 2020), ASASSN-22av (Wu et al. 2024, submitted), and Gaia19bld (Cassan et al., 2022; Rybicki et al., 2022; Bachelet et al., 2022) reached $V\approx 11.5$, $g^{\prime}\approx 12.5$, and $I\approx 9.0$, respectively. The former two were observed using the standard on-axis mode of GRAVITY, whose fringe tracker allow minute-long science exposures (Lacour et al., 2019) that significantly enhance the VLTI sensitivity (Eisenhauer et al., 2023).
@@
Figure 1: Cumulative distribution of the expected $K$-band peak magnitudes of events detected by the OGLE EWS in 2023 (blue line). For comparison, the red line shows the cumulative distribution of the $K$-band magnitudes of the brightest star within $30^{\prime\prime}$ of the event (fringe-tracking star). Only events with $t_{\rm E}\geq 50$ day and an adaptive-optics guide star brighter than $G=14$ are plotted.
@@
The dual-field interferometric observations were designed to alleviate this problem and reach fainter sources (e.g., Shao & Colavita, 1992; Colavita et al., 1999; Delplancke, 2008; Woillez et al., 2014). Still, a nearby bright ($K\!\lesssim\!10$) reference (fringe-tracking) star is needed to stabilize the optical path difference between the telescopes and phase-reference the interferogram of the science target. The angular separation between the science object and the fringe-tracking star must be smaller than $20\!-\!30^{\prime\prime}$, which is set by the atmospheric conditions. In 2019–2022, the GRAVITY instrument was upgraded to enable such dual-field interferometric observations (GRAVITY+ Collaboration et al., 2022), hence the name of this observing mode, GRAVITY Wide.
@@
As soon as ESO offered regular GRAVITY Wide observations since October 2022 (“Period 110”), we initiated a program (PI: Mérand) of interferometric observations of microlensing events. Our primary science goal was to detect and measure precise masses, distances, and transverse velocities of isolated stellar remnants – neutron stars and black holes. Our observations also served as a testbed for verifying the capabilities of the new instrument, and planning, executing, and analyzing the interferometric follow-up observations of a large number of microlensing events.
@@
2 Selection of the Targets for the Interferometric Follow-Up
@@
We began the selection of microlensing targets about 7–10 days before the planned start of each VLTI UT run. The candidates were chosen from publicly available lists of microlensing alerts published by the Optical Gravitational Lensing Experiment (OGLE) Early Warning System [^2] https://ogle.astrouw.edu.pl/ogle4/ews/ews.html (EWS; Udalski, 2003; Udalski et al., 2015), Korean Microlensing Telescope Network (KMTNet) Alert System [^3] https://kmtnet.kasi.re.kr/~ulens/ (Kim et al., 2018), and Microlensing Observations in Astrophysics (MOA) Transient Alerts [^4] https://www.massey.ac.nz/~iabond/moa/alerts/. In addition, we also checked transient alerts published by all-sky surveys, such as Gaia (Hodgkin et al., 2013, 2021) or All Sky Automated Survey for SuperNovae (ASAS-SN) (Shappee et al., 2014).
@@
The selection of targets was based on several scientific and technical criteria. First, we selected events near or past their maximum brightness so that parameters describing their light curves were reasonably well measured. However, the observations had to be secured before the event faded, when the contrast ratio between the microlensed images became too large. We required the contrast ratio to be smaller than 10:1, which corresponds to the maximum lens-source separation (in Einstein radius units) of $u_{\rm max}=1.22$ or, equivalently, the minimum amplitude of $\Delta I_{\rm min}=0.13$.
@@
Because the primary scientific motivation of our project was searching for stellar remnants, which are expected to give rise to long-duration events, we selected events with Einstein timescales longer than $t_{\rm E}=50$ day. However, the nature of the lens cannot be known at the time of selecting the targets. The mass of the lens can be determined only after the interferometric data are combined with the full light curve. That usually means waiting several weeks after the interferometric observations are taken because they are collected close to the maximum magnification.
@@
In the next step, we checked if suitable fringe-tracking and adaptive-optics reference stars are located within $30^{\prime\prime}$ of the event. For possible fringe-tracking stars, we queried the Two Micron All Sky Survey (2MASS) Point Source Catalog (Skrutskie et al., 2006), while for adaptive-optics guide stars – the Gaia Data Release 3 (Gaia Collaboration et al., 2016, 2023). The guide star had to be brighter than $G\approx 14$ for the Multi Application Curvature Adaptive Optics (Arsenault et al., 2003) system, or brighter than $K=8$ for the Coudé Infrared Adaptive Optics (Kendrew et al., 2012) system. Finally, we estimated the expected $K$-band brightness of the event during the planned observations by assuming that the blending parameter in the $K$-band is identical to that in the $I$-band. The baseline $K$-band brightness of the event was taken from the VISTA Variables in the Via Lactea survey (VVV; Minniti et al., 2010), United Kingdom Infra-Red Telescope (UKIRT) Galactic Plane Survey (Lawrence et al., 2007; Lucas et al., 2008), or 2MASS (Skrutskie et al., 2006). We also examined the estimated $K$-band brightness of the source by using the best-fit $I$-band source fluxes as cross checks. For such estimates, we assumed de-reddened color $(I-K)_{0,\rm giant}=1.4$ and $(I-K)_{0,\rm dwarf}=1.0$ for sources roughly classified using the extinction-corrected source magnitudes as giants ($I_{0}<16.5$) and dwarfs ($I_{0}>16.5$), respectively.
@@
Candidate targets were selected independently by two teams (P.M. and S.D.) and subsequently investigated in more detail. In particular, we paid special attention to the microlensing parallax measurements, which are necessary for the lens mass determination. Since the value of the microlensing parallax is inversely proportional to the square root of the lens mass, we required it to be consistent with zero (or close to zero) during the trigger. We also run light curve simulations to ensure that the parallax will be precisely measured (or constrained) by the end of the observing season. We did not consider events with variable source stars because the variability may affect the microlensing parallax measurements.
@@
The blue solid line in Figure 1 shows the cumulative distribution of the expected $K$-band peak magnitudes of events detected by the OGLE’s EWS in 2023. Only events with relatively long timescales ($t_{\rm E}\geq 50$ day) and suitable adaptive-optics guide stars ($G\leq 14$) within $30^{\prime\prime}$ are presented. Only a few events were bright enough ($K<10.5$) for standard GRAVITY on-axis observations. In contrast, the solid red line in Figure 1 shows the distribution of magnitudes of the brightest star within $30^{\prime\prime}$ of the event (which may serve as a fringe-tracking star). Nearly 100 events could have been considered for GRAVITY Wide observations.
@@ 3.1 Photometric Data
OGLE operates the 1.3-m Warsaw Telescope located at Las Campanas Observatory, Chile. The telescope is equipped with a mosaic camera covering a field of view of 1.4 deg2 with the pixel scale of $0.26\,\mathrm{arcsec}\,\mathrm{pixel}^{-1}$. The event has been observed by the OGLE-IV survey since 2010. However, in this paper, we analyze the OGLE data collected from 2016 through 2023 because earlier observations do not contribute to constraining the parameters of the model. Observations were reduced using the OGLE-IV data reduction pipeline (Udalski et al., 2015), which employs a custom implementation of the Difference Image Analysis (DIA) method (Woźniak, 2000).
KMTNet uses three 1.6-m telescopes located at Cerro Tololo Inter-American Observatory (KMTC, Chile), South African Astronomical Observatory (KMTS, South Africa), and Siding Spring Observatory (KMTA, Australia). Each of the KMTNet telescopes is equipped with a camera with 4 deg2 field of view and pixel scale of $0.40\,\mathrm{arcsec}\,\mathrm{pixel}^{-1}$. The analyzed KMTNet observations cover the years 2021–2023. However, because of saturation, we deleted data points near the peak of the event. The KMTNet photometric data were reduced with the tender-loving-care (TLC) DIA-based pipeline (Yang et al., 2024), which was developed from pySIS (Albrow et al., 2009). The vast majority of OGLE and KMTNet images were taken in the $I$-band filter, with additional $V$-band observations to characterize the color of the source star.
Some additional observations were taken in the $R^{\prime}$ band with the 0.18-m Newtonian telescope (CHI-18; Wu et al. 2024) located at the El Sauce Observatory in Chile to cover the peak of the event. The CHI-18 images were reduced with the TLC pipeline. The original observations were taken at a 2-minute cadence. Because the event did not exhibit variability on such short timescales, we binned the CHI-18 data into 1-hour-long bins.

### v2

@@ Abstract
Interferometric observations of gravitational microlensing events offer an opportunity for precise, efficient, and direct mass and distance measurements of lensing objects, especially those of isolated neutron stars and black holes. However, such observations have previously been possible for only a handful of extremely bright events. The recent development of a dual-field interferometer, GRAVITY Wide, has made it possible to reach out to significantly fainter objects and increase the pool of microlensing events amenable to interferometric observations by two orders of magnitude. Here, we present the first successful observation of a microlensing event with GRAVITY Wide and the resolution of microlensed images in the event OGLE-2023-BLG-0061/KMT-2023-BLG-0496. We measure the angular Einstein radius of the lens with subpercent precision, $\theta_{\rm E}=1.280\pm 0.009$ mas. Combined with the microlensing parallax detected from the event light curve, the mass and distance to the lens are found to be $0.472\pm 0.012\,M_{\odot}$ and $1.81\pm 0.05$ kpc, respectively. We present the procedure for the selection of targets for interferometric observations and discuss possible systematic effects affecting GRAVITY Wide data. This detection demonstrates the capabilities of the new instrument, and it opens up completely new possibilities for the follow-up of microlensing events and future routine discoveries of isolated neutron stars and black holes.
@@ 1 Introduction
In his seminal paper on gravitational microlensing, Einstein (1936) realized that the angular separation between the two images of the source star created by a Galactic gravitational lens would be well below the resolution limits of contemporary telescopes. That led him to believe that observing microlensed images would be, in practice, impossible. This notion persisted for many decades until the first optical interferometry facilities were developed.
@@
Although a few tens of thousands of single-lens microlensing events have been discovered so far, only a few have had the angular Einstein radius measured. That was possible for only a small number of events exhibiting finite-source effects (Gould, 1994; Nemiroff & Wickramasinghe, 1994; Witt & Mao, 1994), and the probability of such measurements is biased toward low-mass (planetary-mass) lenses. Another route for $\theta_{\rm E}$ measurements has involved astrometric observations (Hog et al., 1995; Miyamoto & Yoshii, 1995; Walker, 1995), leading to the first detection of an isolated black hole (Sahu et al., 2022; Lam et al., 2022; Mróz et al., 2022; Lam & Lu, 2023). These observations, however, required long-term astrometric monitoring, which is not feasible for a large sample of events. Conversely, interferometric observations are much more efficient, as only one exposure is sufficient to resolve the microlensed images and measure the angular Einstein radius. When combined with the microlensing parallax ($\pi_{\rm E}\equiv\pi_{\rm rel}/\theta_{\rm E}$) measurements from the light curve, they allow us to precisely determine the masses, distances, and transverse velocities of isolated objects, including neutron stars and black holes.
@@
Delplancke et al. (2001) were the first to discuss the prospects of using the European Southern Observatory (ESO) Very Large Telescope Interferometer (VLTI) to study gravitational microlensing events. They estimated that the first-generation VLTI instruments would be able to observe dozens of events every year in the mid 2000s. Yet, the first successful resolution of microlensed images did not materialize until late 2017, when Dong et al. (2019) observed the microlensing event TCP J05074264+2447555 (aka Kojima-1) with the second-generation VLTI instrument GRAVITY (GRAVITY Collaboration et al., 2017).
@@
Why did it take almost two decades to achieve this milestone? The early predictions of VLTI performance underestimated the role and impact of many practical problems encountered during system operations, such as beam stability, vibrations, and air dispersion (e.g., Delplancke, 2008). Moreover, even when these technical problems were solved or mitigated, the performance of a ground-based interferometer is ultimately limited by atmospheric turbulence, which breaks the coherence of the wave fronts arriving at each telescope. Therefore, the observations must be carried out with exposure times ($\sim 1$ ms) shorter than the typical wave-front-coherence timescale (on the order of 20–30 ms) and shorter than the mechanical vibrations timescale, which dominate in the 10–1000 Hz regime. Even for the largest modern 10 m class telescopes, this requirement sets a limiting magnitude of $K\approx 10$ for interferometric observations. Gravitational microlensing events brighter than this limit are exceedingly rare (Figure 1). They occur at most a few times a year, and they exceed the GRAVITY limiting magnitude for a short period of time. The three events with published interferometric observations were unusually bright: TCP J05074264+2447555 (Dong et al., 2019; Zang et al., 2020), ASASSN-22av (Wu et al., 2024a), and Gaia19bld (Cassan et al., 2022; Rybicki et al., 2022; Bachelet et al., 2022) reached $V\approx 11.5$, $g^{\prime}\approx 12.5$, and $I\approx 9.0$, respectively. The former two were observed using the standard on-axis mode of GRAVITY, whose fringe tracker allows minute-long science exposures (Lacour et al., 2019) that significantly enhance the VLTI sensitivity (Eisenhauer et al., 2023).
@@
Figure 1: Cumulative distribution of the expected $K$-band peak magnitudes of events detected by the OGLE EWS in 2023 (blue line). For comparison, the red line shows the cumulative distribution of the $K$-band magnitudes of the brightest star within $30^{\prime\prime}$ of the event (fringe-tracking star). Only events with $t_{\rm E}\geq 50$ days and an adaptive optics guide star brighter than $G=14$ are plotted.
@@
The dual-field interferometric observations were designed to alleviate this problem and reach fainter sources (e.g., Shao & Colavita, 1992; Colavita et al., 1999; Delplancke, 2008; Woillez et al., 2014). In this observing mode, the target and a nearby bright ($K\!\lesssim\!10$) reference (fringe-tracking) star are observed at the same time. The fringe-tracking star is needed to measure the atmospheric turbulence and stabilize the optical path difference between the telescopes, providing a correction to the science target. That enables observation of the target with exposure times longer than the turbulence and vibration timescales, allowing one to reach fainter objects. The angular separation between the science object and the fringe-tracking star must be smaller than $20^{\prime\prime}\!-\!30^{\prime\prime}$, which is set by the atmospheric conditions. In 2019–2022, the GRAVITY instrument was upgraded to enable such dual-field interferometric observations (GRAVITY+ Collaboration et al., 2022), hence the name of this observing mode, GRAVITY Wide.
@@
As soon as ESO offered regular GRAVITY Wide observations from October 2022 (“Period 110”), we initiated a program (PI: A. Mérand) of interferometric observations of microlensing events. Our primary science goal was to detect and measure precise masses, distances, and transverse velocities of isolated stellar remnants—neutron stars and black holes. Our observations also served as a test bed for verifying the capabilities of the new instrument and planning, executing, and analyzing interferometric follow-up observations of a large number of microlensing events.
@@
2 Selection of the Targets for Interferometric Follow-up
@@
We began the selection of microlensing targets about 7–10 days before the planned start of each VLTI UT run. The candidates were chosen from publicly available lists of microlensing alerts published by the Optical Gravitational Lensing Experiment (OGLE) Early Warning System [^3] https://ogle.astrouw.edu.pl/ogle4/ews/ews.html (EWS; Udalski, 2003; Udalski et al., 2015), the Korea Microlensing Telescope Network (KMTNet) Alert System [^4] https://kmtnet.kasi.re.kr/~ulens/ (Kim et al., 2018), and the Microlensing Observations in Astrophysics (MOA) Transient Alerts. [^5] https://www.massey.ac.nz/~iabond/moa/alerts/ In addition, we also checked the transient alerts published by all-sky surveys, such as Gaia (Hodgkin et al., 2013, 2021) and the All Sky Automated Survey for SuperNovae (ASAS-SN; Shappee et al., 2014).
@@
The selection of targets was based on several scientific and technical criteria. First, we selected events near or past their maximum brightness, so that the parameters describing their light curves were reasonably well measured. However, the observations had to be secured before the event faded, before the minor image became too faint, and so the contrast ratio between the microlensed images became too large. We required the contrast ratio to be smaller than 10:1, which was conservatively adopted based on our experience with the GRAVITY data. That contrast ratio corresponds to the maximum lens–source separation (in Einstein radius units) of $u_{\rm max}=1.22$ or, equivalently, the minimum amplitude of $\Delta I_{\rm min}=0.13$.
@@
Because the primary scientific motivation of our project was searching for stellar remnants, which are expected to give rise to long-duration events, we selected events with Einstein timescales longer than $t_{\rm E}=50$ days. However, the nature of the lens cannot be known at the time of selecting the targets. The mass of the lens can be determined only after the interferometric data are combined with the full light curve. That usually means waiting several weeks after the interferometric observations are taken, because they are collected close to the maximum magnification.
@@
In the next step, we checked if suitable fringe-tracking and adaptive optics reference stars were located within $30^{\prime\prime}$ of the event. For possible fringe-tracking stars, we queried the Two Micron All Sky Survey (2MASS) Point Source Catalog (Skrutskie et al., 2006), while for adaptive optics guide stars, we queried Gaia Data Release 3 (Gaia Collaboration et al., 2016, 2023). The guide star had to be brighter than $G\approx 14$ for the Multi Application Curvature Adaptive Optics (Arsenault et al., 2003) system or brighter than $K=8$ for the Coudé Infrared Adaptive Optics (Kendrew et al., 2012) system. Finally, we estimated the expected $K$-band brightness of the event during the planned observations, by assuming that the blending parameter in the $K$ band was identical to that in the $I$ band. The baseline $K$-band brightness of the event was taken from the VISTA Variables in the Via Lactea (VVV) survey (Minniti et al., 2010), the United Kingdom Infra-Red Telescope (UKIRT) Galactic Plane Survey (Lawrence et al., 2007; Lucas et al., 2008), or 2MASS (Skrutskie et al., 2006). We also examined the estimated $K$-band brightness of the source by using the best-fit $I$-band source fluxes as cross-checks. For such estimates, we assumed dereddened color $(I-K)_{0,\rm giant}=1.4$ and $(I-K)_{0,\rm dwarf}=1.0$ for sources roughly classified using the extinction-corrected source magnitudes as giants ($I_{0}<16.5$) and dwarfs ($I_{0}>16.5$), respectively.
@@
Candidate targets were selected independently by two teams (P.M. and S.D.) and subsequently investigated in more detail. In particular, we paid special attention to the microlensing parallax measurements, which are necessary for the lens mass determination. Because the value of the microlensing parallax is inversely proportional to the square root of the lens mass, we required it to be consistent with zero (or close to zero) during the trigger. We also ran light-curve simulations to ensure that the parallax would be precisely measured (or constrained) by the end of the observing season. We did not consider events with variable source stars, because the variability may affect the microlensing parallax measurements.
@@
The blue solid line in Figure 1 shows the cumulative distribution of the expected $K$-band peak magnitudes of the events detected by the OGLE EWS in 2023. Only events with relatively long timescales ($t_{\rm E}\geq 50$ days) and suitable adaptive optics guide stars ($G\leq 14$) within $30^{\prime\prime}$ are presented. Only a few events were bright enough ($K<10.5$) for standard GRAVITY on-axis observations. In contrast, the solid red line in Figure 1 shows the distribution of the $K$-band magnitudes of the brightest star within $30^{\prime\prime}$ of the event (which may serve as a fringe-tracking star). Nearly 100 events could have been considered for GRAVITY Wide observations.
@@ 3.1 Photometric Data
OGLE operates the 1.3 m Warsaw Telescope located at Las Campanas Observatory, Chile. The telescope is equipped with a mosaic camera covering a field of view of 1.4 deg2 with a pixel scale of $0.26\,\mathrm{arcsec}\,\mathrm{pixel}^{-1}$. The event has been observed by the OGLE-IV survey since 2010. However, in this paper, we analyze the OGLE data collected from 2016 through 2023, because earlier observations do not contribute to constraining the parameters of the model. Observations were reduced using the OGLE-IV data reduction pipeline (Udalski et al., 2015), which employs a custom implementation of the Difference Image Analysis (DIA) method (Woźniak, 2000).
KMTNet uses three 1.6 m telescopes located at the Cerro Tololo Inter-American Observatory (KMTC; Chile), the South African Astronomical Observatory (KMTS; South Africa), and the Siding Spring Observatory (KMTA; Australia). Each of the KMTNet telescopes is equipped with a camera with a 4 deg2 field of view and a pixel scale of $0.40\,\mathrm{arcsec}\,\mathrm{pixel}^{-1}$. The analyzed KMTNet observations cover the years 2021–2023. However, because of saturation, we deleted data points near the peak of the event. The KMTNet photometric data were reduced with the tender-loving-care (TLC) DIA-based pipeline (Yang et al., 2024), which was developed from pySIS (Albrow et al., 2009). The vast majority of OGLE and KMTNet images were taken in the $I$-band filter, with additional $V$-band observations to characterize the color of the source star.
Some additional observations were taken in the $R^{\prime}$ band with the 0.18 m Newtonian telescope (CHI-18; Wu et al. 2024b), located at the El Sauce Observatory in Chile, to cover the peak of the event. The CHI-18 images were reduced with the TLC pipeline. The original observations were taken at a 2 minutes cadence. Because the event did not exhibit variability on such short timescales, we binned the CHI-18 data into 1 hr long bins.
@@
Figure 2: UV plane coverage of epoch 1 (left panel) and epoch 2 (right panel) GRAVITY observations of OGLE-2023-BLG-0061/KMT-2023-BLG-0496.